Jonah invests R4500 in a savings account. The interest rate for the first 4 years is 9 % per year, compounded monthly, thereafter the interest rate changes to 10% per year, compounded half yearly, for the next 5 years. Determine the amount that Jonah will have in his savings account at the end of the term.


1
Expert's answer
2021-09-17T03:15:39-0400

for the first 4 years,

"A=P(1+r\/100)^n"

"r=9/12=0.75\% , n=4\times 12=36, P=R4500"

"A=45000(1+0.75/100)^{36}=4500\times 1.0075^{36}=R5888.9"

for the next 5 years

"r=10/2=5\%, n=5\times 2=10, P=R5888.9"

"A=5888.9(1+5/100)^{10}=5888.9\times 1.05^{10}=R9592.40"